OpenStudy (anonymous):

Will Medal What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (gabebae):

\[40+35= 5x+35\]\[75= 5x+ 35\]\[40= 5x\]\[8= x\]

OpenStudy (anonymous):

@GabeBae is awesome

OpenStudy (gabebae):

lel